1. A Van de Graaff generator has a metal sphere on the top with a radius of 15.0 cm. It is charged to the point where the total charge on the surface of the sphere is +3.50 µC. (a) What sort of particles have (probably) been added to or removed from the surface of the Van de Graaff generator? How many of them have been added or removed? (b) A small metal ball (1.00 cm radius) mounted on an insulating stand is placed a large distance from the Van de Graaff generator. It is initially neutral. It is then connected to the Van de Graaff generator with a long, thin wire. What particles flow and which way do they flow through the wire (i.e. do they flow from the generator to the ball, or the other way)? (c) The small metal ball was kept a long distance from the Van de Graaff generator because this makes things simpler. How does this simplify the situation? (d) On a future assignment (probably Assignment #4) you will show that after the wire is connected, the small ball gains a charge of +219 nC. How many charged particles were transferred between the sphere and the Van de Graaff generator? (e) What is the charge on the Van de Graaff generator now? (f) Suppose that by a “long distance” we mean 2.00 m (this is long enough for our purposes). In other words, the center of the metal ball is 2.00 m away from the center of the Van de Graaff generator. How strong is the force that the Van de Graaff generator exerts on the sphere? Would this likely be noticeable? (g) In part 1f you had to make an approximation (did you realize you were making an approximation?). If the ball was not a “long distance” from the Van de Graaff generator would have been able to make this approximation? Would you be able to do the calculation of the forces that the objects exert on each other? (Hint: think about part 1c.) (h) Ions are constantly generated, naturally, in air. This often happens because of cosmic rays or ultraviolet light interacting with the air. Suppose that an N+ 2 ion is produced near to the metal ball and the Van de Graaff generator. The Van de Graaff generator is at the origin, the metal ball is at 2.00ˆi m and the ion is at 2.00ˆi + 0.300ˆj m. Find the acceleration (vector) of the ion
Answer #1

a)

Since the sphere gains positive charge , electrons have been removed from it .

Q = total change gained = 3.50 x 10-6 C

q = change on single charge carrier = 1.6 x 10-19 C

n = number of carriers

using the equation

Q = n q

3.50 x 10-6 = n (1.6 x 10-19)

n = 2.1875 x 1013

b)

the metal sphere has deficiency of electrons , hence the electrons from small metal ball travels to the metal sphere until the electric potential on the two spheres becomes equal.

c)

placing the ball large distance away is necessary to avoid charging due to induction or polarization.

d)

Q = charge transferred = 219 nC = 219 x 10-9 C

n = number of particles

q = charge on single particle = 1.6 x 10-19 C

using the equation

Q = n q

219 x 10-9 = n (1.6 x 10-19)

n = 1.36875 x 1012